solve using laplace please.

xy'' + (1-x)y' + 2y = 0 , y(0) = 1 , y'(0) = -2
Answer #1

Given equation is. xy+(1-2) Ytayo yeo) = 1, y60) = -2. 3 -5 SCS-1) ay + y - xy + 2y = 6 we know that Llyn = s²yes) - syo)- yCS) [-s²+s] + y(s) [-s +3] +1-1 so. yles) (-5°+s] + y) (-s+] = 0 yyles) ( S2-s]+ yes) [S-3) = 0. => Y [SP-] - Y(S) [3-5).- lny = ln s 3 + ln (S-1)2 + lnc any en (c(s.1²) .: yes) - co Yes) = ( (5²_25+1) 53 yes) =[s B transform on both sides laplac